Marko invested $300 000 in stocks, bonds and a savings account at the rates shown on the right. (stocks=15%, bonds=10%, savings=4%). He invested 4 times as much in stocks as he invested in the savings account. After one year, he earned $35 600 in interest. How much money did he put into each type of investment?
Answer by ankor@dixie-net.com(22740)

Marko invested $300 000 in stocks, bonds and a savings account at the rates shown on the right.
(stocks=15%, bonds=10%, savings=4%).
He invested 4 times as much in stocks as he invested in the savings account. After one year, he earned $35 600 in interest.
How much money did he put into each type of investment?
let s = amt invested in stocks
let b = amt invested in bonds
let a = amt in savings
:
s + b + a = 300000
and
.15s + .10b + .04a =35600
"He invested 4 times as much in stocks as he invested in the savings"
s = 4a
:
replace s with 4a in both original equation
4a + b + a = 30000
5a + b = 300000
and
.15(4a) + .10b + .04a = 35600
.60a + .10b + .04a = 35600
.64a + .10b = 35600
Multiply by 10 and subtract the 1st two unknown equation
6.4a + b = 356000
5a + b = 300000
------------------subtraction eliminates b, find a
1.4a = 56000
a = 56000/1.4
a = $40,000 in savings
then
s = 4(40000)
s = $160,000 in stocks
then
300000 - 40000 - 160000 = $100,000 in bonds
:
:
:
Check this
.15(160000) + .10(100000) + .04(40000) = 35600